find two positive numbers whose squares have the sum of 85 and difference of 13
Answer by oberobic(2304)

x^2 - y^2 = 13
x^2 + y^2 = 85
--------------
2x^2 = 98
x^2 = 49
x = +/- 7, but we're told to pick the positive value, so
x = 7
substitute to find 'y'
(7^2) + y^2 = 85
y^2 = 85-49
y^2 = 36
y = 6
check these values in the other equation
7^2 -6^2 = 13 ??
49-36=13
correct.
Answer: The two numbers are 6 and 7.
Done.